Sundance Solar Company operates two factories. The company applies factory overhead to jobs on the basis of machine hours inRequired: a. Determine the factory overhead rate for Factory 1. b. Determine the factory overhead rate for Factory 2. C. JourCHART OF ACCOUNTS Sundance Solar Company General Ledger ASSETS REVENUE 110 Cash 410 Sales 121 Accounts Receivable 610 Interes133 Factory Overhead 134 Finished Goods 531 Selling Expenses 532 Insurance Expense 533 Utilities Expense 141 Supplies 142 PreLIABILITIES 210 Accounts Payable 221 Utilities Payable 231 Notes Payable 236 Interest Payable 241 Lease Payable 251 Wages Pay351 Dividends 390 Income Summarya. Determine the factory overhead rate for Factory 1. per machine hour b. Determine the factory overhead rate for Factory 2.C. Journalize the Mar. 31 entry to apply factory overhead to production in Factory 1. Refer to the Chart of Accounts for exacNow journalize the second Mar. 31 entry to apply factory overhead to production in Factory 2. Refer to the Chart of Accountsd. Determine the balances of the factory overhead accounts for each factory as of March 31, and indicate whether the amounts
Answer #1
a
Factory 1 30 per machine hour =1504500/50150
b
Factory 2 36 per direct labor hour =914400/25400
c
Date Description Post. Ref. Debit Credit
March 31 Work in Process 123000 =4100*30
     Factory Overhead 123000
March 31 Work in Process 105480 =2930*36
     Factory Overhead 105480
d
Factory 1 1420 Credit Overapplied =123000-121580
Factory 2 1720 Debit Underapplied =107200-105480